David Ortiz Motors has a target capital structure of 45% debt and 55% equity. The yield to maturity on the company's outstanding bonds is 10%, and the company's tax rate is 40%. Ortiz's CFO has calculated the company's WACC as 8.97%. What is the company's cost of equity capital? Round your answer to two decimal places.
Answer #1

first let us know the after tax of debt => yield to maturity *(1- tax rate)

=>10%*(1-0.40)

=>6%.

now,

WACC = (weight of debt* after tax cost of debt) + (weight of equity * cost of equity)

=> (0.45*6%) + (0.55* cost of equity) = 8.97%

=> (2.70%) + (0.55* cost of equity) = 8.97%

=>0.55 * cost of equity = 8.97-2.70

=>0.55* cost of equity = 6.27

=>cost of equity = 6.27 /0.55.

=>cost of equity =11.40%.